Opened 10 years ago

Closed 4 years ago

#324 closed flight processing (fixed)

GB07/12, flight day 082b/2010, Debden

Reported by: benj Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2010/flight_data/uk/GB07_12-2010_082b_Debden

Data arrived from ARSF via network transfer on 25th March 2010.

Scientific objective: Monitoring of peat fires and quantification of pyrogenic carbon conversion

Priority: a4l

PI: Martin Wooster

Note original PI on application was Andrew Wilson, but since he's retired effective PI is Martin Wooster.

Sensors:

  • Eagle (07/09/2010) (reprocessed 23/04/2012)
  • Hawk (20/12/2010) (reprocessed 23/04/2012)
  • Leica LIDAR (19/07/2010)
  • RCD (23/07/2010)

Change History (46)

comment:1 Changed 10 years ago by benj

  • Description modified (diff)
  • Summary changed from GB07/12, flight day 082b/2010, Debden burn to GB07/12, flight day 082b/2010, Debden

comment:2 Changed 10 years ago by benj

  • Milestone set to 2010 data processing completion

comment:3 Changed 10 years ago by mark1

removed rcd logs that related to flight 082a Debden

comment:4 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Navigation data is already processed for this project - possibly came from 082a?

comment:5 Changed 10 years ago by benj

I think Chris auto-Applanixed this one so it was available to be sent to Pol for use with the TABI (which hasn't happened yet).

comment:6 Changed 10 years ago by pmlvis10

Navigation processing done.

comment:7 Changed 10 years ago by pmlvis10

*IPAS navigation processing done!

comment:8 Changed 10 years ago by knpa

started LIDAR processing

comment:9 Changed 10 years ago by knpa

Lines processed. Currently doing QC.

Can find no roll for this project and no major vertical or horizontal discrepancies have been found.

comment:10 Changed 10 years ago by knpa

Classification done.

comment:11 Changed 10 years ago by knpa

LIDAR delivery directory created and populated. Need typed up logsheet to create Readme, Ops say it should be on the way soon.

comment:12 Changed 10 years ago by mark1

Typed logsheet has been received from Ops and added to admin dir.

comment:13 Changed 10 years ago by knpa

LIDAR delivery awaiting delivery check.

Started Hyperspectral processing.

comment:14 Changed 10 years ago by mark1

LIDAR delivery checked: - need to:

Update data-quality-report to 2010 one (get from emca if not in ~arsf/doc).
Use the typed (not scanned) logsheet - rename the lines so that they dont all read "1 fire" or "4 fire"
In Readme:

Update table to reflect new names
Use full filename or name from logsheets in per-line stats and elevation comparisons.
Site location should be Debden


comment:15 Changed 10 years ago by knpa

Eagle flightline raw file names are a bit muddled.

From the hdr times, it looks like:

-1 = line 1
.1-1 & .1.1-1 = line 2
-2 = line 3
-3 = line 4
...etc

Renaming files to match correct flightline number. Also adding VNIR prefix (and SWIR to hawk).

comment:16 Changed 10 years ago by knpa

Updated logsheet and Readme according to Mark1's above instructions.

Flight names now given a suffix according to their position on logsheet, producing 1 fire (a-e) and 4 fire (a-f).

However, latest calibration report is not ready. Unable to delivery lidar until this is finished.

comment:17 Changed 10 years ago by knpa

Renamed eagle and hawk files.
Now have VNIR and SWIR prefixes respectively.
Eagle now matches flightlines, with -2.1 and -2.2 for the line that is in two parts.

comment:18 Changed 10 years ago by knpa

Told by emca that the lidar calibration file (day 060) was changed shortly after it was used to process this project. The range corrections in the file were originally wrong and were changed by ~2m.
As a result this project is ~2m lower than it should be. Having to reprocess with correct range cal settings...

comment:19 Changed 10 years ago by knpa

Flightlines reprocessed. Generated new DEM and updated data quality report.

Requested Mark1 to recheck before sending out.

comment:20 Changed 10 years ago by mark1

LIDAR delivery checked and looks OK, renamed DEM to delivery format.

comment:21 Changed 10 years ago by knpa

  • Description modified (diff)

LIDAR delivered today (19/07/2010)

comment:22 Changed 10 years ago by knpa

  • Description modified (diff)

RCD data for this flight too!

comment:23 Changed 10 years ago by mark1

Camera images have been processed. Removed 1-13 since these do not cover the survey site.

Delivery directory has been created and is ready for checking at:
~airborne/workspace/GB07_12-2010_082b_Debden/delivery/20100722/GB07_12

comment:24 Changed 10 years ago by emca

Delivery checked photographs. All looks good. Ready to deliver.

comment:25 Changed 10 years ago by mark1

  • Description modified (diff)

Photographs dispatched to PI (23/07/2010)

comment:26 Changed 10 years ago by knpa

Renamed eagle raw files again so they match those stated on logsheet (i.e. line 2 is now comprised of raw files 2 & 3 rather than 2.1 and 2.2).

comment:27 Changed 10 years ago by knpa

Eagle was ran full spectral resolution (505 bands) and looks noisy as a result.

comment:28 Changed 10 years ago by knpa

Eagle is missing many frames, probably due to running at full spectral res.

comment:29 Changed 10 years ago by knpa

SCTs

eagle hawk
1 -0.04 0.02
2 0.00 0.03
3 0.02 0.02
4 -0.01 0.02
5 0.00 0.04
6 0.05 0.01
7 0.01 0.02
8 0.00 0.02
9 -0.03 1.01 (sync)
10 0.01 0.02
11 0.00 0.02

comment:30 Changed 10 years ago by knpa

and eagle 12: 0.00

comment:31 Changed 10 years ago by knpa

Delivery created, awaiting delivery check.

comment:32 Changed 10 years ago by jb362

Starting hyperspectral delivery check

comment:33 Changed 10 years ago by jb362

Delivery check complete. Everything fine.

comment:34 Changed 10 years ago by knpa

e-mailed Martin Wooster to ask if he would like an early delivery on Eagle.
Auto-reply says he is away until 20th Auguest so this project will be on hold until then (or until Hawk issues are resolved).

comment:35 Changed 10 years ago by mark1

  • Description modified (diff)

Eagle data dispatched to PI on 07/09/2010

comment:36 Changed 9 years ago by anch

Hawk delivery check complete:

Everything looks fine

comment:37 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data sent 08/12/2010

comment:38 Changed 9 years ago by emca

Hawk data reprocessed and ready for delivery check

comment:39 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data re-delivered 20/12/2010

comment:40 Changed 8 years ago by emca

Reprocessed hyperspectral data using apl.

Data sent to PI 23/04/2012

comment:41 Changed 8 years ago by knpa

Have r-synced the workspace version (non-reprocessed).

comment:42 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving

Beginning archiving.

comment:43 Changed 7 years ago by knpa

Made available to NDEODC.

comment:44 Changed 7 years ago by knpa

Backed up by NEODC.

Moved to archive disk.

comment:45 Changed 6 years ago by knpa

  • Description modified (diff)

comment:46 Changed 4 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/GB07_12/GB07_12-2010_082b_Debden). Marking as closed.

Note: See TracTickets for help on using tickets.